# On the solution of the Collatz problem

### v26

@@ 3.3. Properties of odd numbers $n_{0}(m,t)$ in a range $(N,\rho N]$
In the case of $r_{1}=2r$ where $r>0$, the odd number $n=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.22, when one of the two numbers
$13\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1
returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$. So when one of the two odd numbers $n$ and $13^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1 another also returns to 1. Let $n_{r}=3^{r}n_{i_{1}}(m_{1},t_{1})$ where one of the two odd numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.18 the odd number $n_{r}$ returns to 1, then the even number $13n_{r}+1=3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, and by Lemma 3.22 the odd number $n=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1.
In the case of $r_{1}=2r+1$ where $r>0$, the odd number $n=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.23, when one of the two odd numbers
$13\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1
returns to 1 another also returns to 1. So when one of the two odd numbers $n$ and $13^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1 another also returns to 1. Let $m$ and $t$ be determined by
@@
where $0\leq i<m$ and one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.17, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m_{1},t_{1})$ returns to 1 another also returns to 1, and when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.18 $S_{c}(m-1,t)$ returns to 1, and by Lemma 3.17 $S_{c}(m,t)$ returns to 1, i.e., the odd number $13^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, so that by Lemma 3.23 the odd number $n=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1.
@@
Given an odd number $n=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})+1$ where $r_{1}>0$, $0\leq i_{1}\leq m_{1}$, and $n_{i_{1}}(m_{1},t_{1})$ is odd and not dividable by 3, when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
By Lemma 3.5 a half of positive integers $r_{1}$ satisfy $r_{1}=1$ and a quarter of positive integers $r_{1}$ satisfy $r_{1}=2$.
In the case of $r_{1}=1$, the odd number $n=2n_{i_{1}}(m_{1},t_{1})+1$. Let $m=m_{1}+1$ and $t=t_{1}$. Then by Lemma 3.17, when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
where $0\leq i<m$ and one of the two odd numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two Collatz’s Sequences $S_{c}(m-1,t)$ and $S_{c}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.17, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
In the case of $r_{1}=2r+2$ where $r>0$, there is $n_{1}=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.22, when one of the two numbers
$13\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1
returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$. So when one of the two odd numbers $n_{1}$ and $13^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1 another also returns to 1. Let $n_{r}=3^{r}n_{i_{1}}(m_{1},t_{1})$ where one of the two odd numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.18 the odd number $n_{r}$ returns to 1, then the even number $13n_{r}+1=3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, and by Lemma 3.22 the odd number $n_{1}=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that the odd number $n$ returns to 1.
In the case of $r_{1}=2r+3$ where $r>0$, there is $n_{1}=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.23, when one of the two odd numbers
$13\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1
returns to 1 another also returns to 1. So when one of the two odd numbers $n_{1}$ and $13^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1 another also returns to 1. Let $m$ and $t$ be determined by
@@
where $0\leq i<m$ and one of the two odd numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.17, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.18 $S_{c}(m-1,t)$ returns to 1, and by Lemma 3.17 $S_{c}(m,t)$ returns to 1, i.e., the odd number $13^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, then by Lemma 3.23 the odd number $n_{1}=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that the odd number $n$ returns to 1.
@@
By Lemma 3.5 a half of positive integers $r_{1}$ satisfy $r_{1}=1$ and a quarter of positive integers $r_{1}$ satisfy $r_{1}=2$. Let the notation $P_{r}$ denote the process of mutual references or nested recursions to lemmas 3.17-3.20.
In the proof of Lemma 3.17, a half of positive integers $r_{1}$ satisfy $r_{1}=1$ and directly make Lemma 3.17 true, another half of positive integers $r_{1}$ continue the process $P_{r}$.
In the proof of Lemma 3.20, a quarter of positive integers $r_{1}$ satisfy $r_{1}=2$ and directly make Lemma 3.20 true, and a half of positive integers $r_{1}$ satisfy $r_{1}=1$ and make the reference to Lemma 3.17, which is equivalent to that among these positive integers $r_{1}=1$ a half of them directly make Lemma 3.20 true and another half of them continue the process $P_{r}$. So it is equivalent to that a half of positive integers $r_{1}$ satisfy $r_{1}=1,2$ and directly make Lemma 3.20 true and another half of positive integers $r_{1}$ satisfy $r_{1}\neq 2$ and continue the process $P_{r}$.
@@
When one of the two numbers
$13\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1
@@
So when one of the two numbers
$13\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1
@@
When one of the two odd numbers
$13\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1
@@
So when one of the two odd numbers
$13\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1

### v32

@@ 3.3. Properties of odd numbers $n_{0}(m,t)$ in a range $(N,\rho N]$
In the case of $r_{1}=2r$ where $r>0$, the odd number $n=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.27, when one of two numbers
$3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$
returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$. So when one of two numbers $n$ and $3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1 another also returns to 1. Let $n_{r}=3^{r}n_{i_{1}}(m_{1},t_{1})$ where one of two numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.23 the odd number $n_{r}$ returns to 1, then the even number $3n_{r}+1=3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, and by Lemma 3.27 the odd number $n=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1.
In the case of $r_{1}=2r+1$ where $r>0$, the odd number $n=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.28, when one of two numbers
$3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$
returns to 1 another also returns to 1. So when one of two numbers $n$ and $3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1 another also returns to 1. Let $m$ and $t$ be determined by
@@
where $0\leq i<m$ and one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.22, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m_{1},t_{1})$ returns to 1 another also returns to 1, and when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.23 $S_{c}(m-1,t)$ returns to 1, and by Lemma 3.22 $S_{c}(m,t)$ returns to 1, i.e., the odd number $3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that by Lemma 3.28 the odd number $n=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1.
@@
Given an odd number $n=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})+1$ where $r_{1}>0$, $0\leq i_{1}\leq m_{1}$, and $n_{i_{1}}(m_{1},t_{1})$ is odd and not dividable by 3, when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
By Lemma 3.5 one half of positive integers $r_{1}$ satisfy $r_{1}=1$ and one quarter of positive integers $r_{1}$ satisfy $r_{1}=2$.
In the case of $r_{1}=1$, the odd number $n=2n_{i_{1}}(m_{1},t_{1})+1$. Let $m=m_{1}+1$ and $t=t_{1}$. Then by Lemma 3.22, when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
where $0\leq i<m$ and one of two numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of the two Collatz’s Sequences $S_{c}(m-1,t)$ and $S_{c}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.22, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
In the case of $r_{1}=2r+2$ where $r>0$, there is $n_{1}=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.27, when one of two numbers
$3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$
returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$. So when one of two numbers $n_{1}$ and $3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1 another also returns to 1. Let $n_{r}=3^{r}n_{i_{1}}(m_{1},t_{1})$ where one of two numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n_{r}$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.23 the odd number $n_{r}$ returns to 1, then the even number $3n_{r}+1=3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, and by Lemma 3.27 the odd number $n_{1}=3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that the odd number $n$ returns to 1.
In the case of $r_{1}=2r+3$ where $r>0$, there is $n_{1}=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$. By Lemma 3.28, when one of two numbers
$3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$
returns to 1 another also returns to 1. So when one of two numbers $n_{1}$ and $3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1 another also returns to 1. Let $m$ and $t$ be determined by
@@
where $0\leq i<m$ and one of two numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.22, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, by Lemma 3.23 $S_{c}(m-1,t)$ returns to 1, and by Lemma 3.22 $S_{c}(m,t)$ returns to 1, i.e., the odd number $3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, then by Lemma 3.28 the odd number $n_{1}=3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that the odd number $n$ returns to 1.
@@
By Lemma 3.5 one half of positive integers $r_{1}$ satisfy $r_{1}=1$ and one quarter of positive integers $r_{1}$ satisfy $r_{1}=2$. Let the notation $P_{r}$ denote the process of mutual references or nested recursions to lemmas 3.22-3.25.
In the proof of Lemma 3.22, one half of positive integers $r_{1}$ satisfy $r_{1}=1$ and directly make Lemma 3.22 true, another half of positive integers $r_{1}$ continue the process $P_{r}$.
In the proof of Lemma 3.25, one quarter of positive integers $r_{1}$ satisfy $r_{1}=2$ and directly make Lemma 3.25 true, and one half of positive integers $r_{1}$ satisfy $r_{1}=1$ and make the reference to Lemma 3.22, which is equivalent to that among these positive integers $r_{1}=1$ one half of them directly make Lemma 3.25 true and another half of them continue the process $P_{r}$. So it is equivalent to that one half of positive integers $r_{1}$ satisfy $r_{1}=1,2$ and directly make Lemma 3.25 true and another half of positive integers $r_{1}$ satisfy $r_{1}\neq 2$ and continue the process $P_{r}$.
@@
When one of two numbers
$3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$
@@
So when one of two numbers
$3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$
@@
When one of two numbers
$3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$
@@
So when one of two numbers
$3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$